OpenStudy (anonymous):

the position of a moving particle is given s(t)=2t^3-21t^2+60t where distance is measured in ft & time in secs. find velocity of particle at t=1

OpenStudy (anonymous):

s(t)=2t^3-21t^2+60t v(t)=d(s(t))/dt (v=velocity) =6t^2-42t+60 by puting the value of t we get v=6-42+60 =24 ft/sec